Costs, income and profit maximization

Instructions

You must calculate with the information provided everything that is in blue and yellow, once the calculations made you must answer the questions below the table. I remind you that the mathematical procedure must be present in the equations, otherwise the exercise is Invalid.

Evaluation: Each question has a value of 8 points and making the table has a value of 4 points, you must be aware of putting the numbers in the table correctly, because if the number is wrongly placed in the table the error will drag it even the formulas.

Suppose you are the owner of a telephone factory, the capacity of the plant is 60 thousand units, the sale price is $ 50.00. Variable costs are represented as follows: Direct Materials 20, Labor 15 and Overhead 7, the Total Fixed Cost is $ 480,000.

Break Even Table:

Break Even Chart

Production

10,000

20,000

30,000

40,000

50,000

60,000

Total Revenue (P x Q)

Cost

Total Direct Material (DM x Q)

Total Direct labor (DL x Q)

Total Overhead (O x Q)

Total Variable Cost

Total Fixed Cost (TFC)

Total Cost (TVC + TFC)

Gain or Loss (Total revenue- Total Cost)

1) How many units must the company produce and sell to avoid losses or

       Profits ?.

2) What is the number of units that the company must produce or sell in order to cover all its costs and also obtain the expected profit? Suppose your company wants to make a profit of $ 90,000.

    3) How many units must it produce or sell in order to cover its total costs, pay taxes and obtain an expected profit?

4) What is the income that the company must obtain by selling its production to avoid having a loss or gain?

5) If your company has an interest in covering the fixed costs, the variable costs plus the expected profit, what is the income that it must have to cover all this?
Answer #1

330 Sous Sox) TA 3oC 1966 SA M Sore (72) Dur Brot 1300 o 485D KLO U Bass woooo g 3 Stup 8100 X 150 Blue S End Lot BU YOLD DC

1. The company should produce all 60000 units to be at break even point that is the situation where there is no profit and no loss to the company.

2. Number of units to be produced to achieve desired profit =(desired profit+ fixed cost)/contribution per unit.

(90000+480000)/8= 71250 units.

Contribution per units can be calculated from table =80000/10000=8.

4.To cover all costs at break even point the company should earn a revenue of 3000000.

5. Income= 480000+90000+42x

Where x= no. Of units produced.

42 as 20+15+7

480000= fixed cost

90000= desired profit